40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12571     Entry time: Wed Oct 19 16:41:55 2016     In reply to: 12568
Author: gautam 
Type: Update 
Category: General 
Subject: Heavy doors back on 

[ericq, lydia, steve, gautam]

  • We aligned the arms, and centered the in-air AS beam onto the PDs and camera
  • Misaligned the ITMs in a controlled ramp, observed ASDC level, didn't see any strange features
  • We can misalign the ITMs by +/- 100urad in yaw and not see any change in the ASDC level (i.e. no clipping). We think this is reasonable and it is unlikely that we will have to deal with such large misalignments. We also scanned a much larger range of ITM misalignments (approximately +/-1mrad), and saw no strange features in the ASDC levels as was noted in this elog - we used both the signal from the AS110 PD which had better SNR and also the AS55 PD. We take this to be a good sign, and will conduct further diagnostics once we are back at high power.
  • Opened up all light doors, checked centering on all 6 OM mirrors again, these were deemed to be satisfactory 
  • To solve the green scattering issue, we installed a 1in wide glass piece (~7inches tall) mounted on the edge of the OMC table to catch the reflection off the window (see Attachment #1) - this catches most of the ghost beams on the PSL table, there is one that remains directly above the beam which originates at the periscope in the BS/PRM chamber (see Attachment #2) but we decided to deal with this ghost on the PSL table rather than fiddle around in the vacuum and possibly make something else worse
    Link to IMG_2332.JPG
    Link to IMG_2364.JPG
  • Re-aligned arms, ran the dither, and then aligned the PRM and SRM - we saw nice round DRMI flashes on the cameras
  • Took lots of pictures in the chamber, put heavy doors back on. Test mass Oplev spots looked reasonably well centered, I re-centerd PRM and SRM spots in their aligned states, and then misaligned both
  • The window from the OMC chamber to the AS table looked clean enough to not warrant a cleaning..
  • PSL shutter is closed for now. I will check beam alignment, center Oplevs, and realign the green in the evening. Plan is to pump down first thing tomorrow morning

AS beam on OM1

Link to IMG_2337.JPG

AS beam on OM2

AS beam on OM3

AS beam on OM4

 
AS beam on OM6

I didn't manage to get a picture of the beam on OM5 because it is difficult to hold a card in front of it and simultaneously take a photo, but I did verify the centering...

It remains to update the CAD diagram to reflect the new AS beam path - there are also a number of optics/other in-vacuum pieces I noticed in the BS/PRM and OMC chambers which are not in the drawings, but I should have enough photos handy to fix this.  

Here is the link to the Picasa album with a bunch of photos from the OMC, BS/PRM and ITMY chambers prior to putting the heavy doors back on...


SRM satellite box has been removed for diagnostics by Rana. I centered the SRM Oplev prior to removing this, and I also turned off the watchdog and set the OSEM bias voltages to 0 before pulling the box out (the PIT and YAW bias values in the save files were accurate). Other Oplevs were centered after dither-aligning the arms (see Attachment #8, ignore SRM). Green was aligned to the arms in order to maximize green transmission (GTRX ~0.45, GTRY ~0.5, but transmission isn't centered on cameras).

I don't think I have missed out on any further checks, so unless anyone thinks otherwise, I think we are ready for Steve to start the pumpdown tomorrow morning.

Attachment 1: IMG_2332.JPG  2.383 MB  Uploaded Wed Oct 19 20:13:20 2016  | Hide | Hide all
IMG_2332.JPG
Attachment 2: IMG_2364.JPG  2.092 MB  Uploaded Wed Oct 19 20:18:13 2016  | Hide | Hide all
IMG_2364.JPG
Attachment 3: IMG_2337.JPG  2.278 MB  Uploaded Wed Oct 19 20:20:25 2016  | Hide | Hide all
IMG_2337.JPG
Attachment 4: IMG_2338.JPG  2.277 MB  Uploaded Wed Oct 19 20:20:45 2016  | Hide | Hide all
IMG_2338.JPG
Attachment 5: IMG_2356.JPG  1.835 MB  Uploaded Wed Oct 19 20:22:04 2016  | Hide | Hide all
IMG_2356.JPG
Attachment 6: IMG_2357.JPG  1.810 MB  Uploaded Wed Oct 19 20:22:17 2016  | Hide | Hide all
IMG_2357.JPG
Attachment 7: IMG_2335.JPG  2.532 MB  Uploaded Wed Oct 19 20:22:48 2016  | Hide | Hide all
IMG_2335.JPG
Attachment 8: Oplevs_19Oct2016.png  17 kB  Uploaded Wed Oct 19 20:35:33 2016  | Hide | Hide all
Oplevs_19Oct2016.png
ELOG V3.1.3-